OpenStudy (anonymous):

use mclaurin formula f(x)= tan x up to the term x^5

OpenStudy (amistre64):

i never really knew the difference between taylor and mcclaurin ... one has to do with 0 and the other with a

OpenStudy (amistre64):

\[P(x)=f(a)+\frac{f'(a)}{1!}(x-a)+\frac{f''(a)}{2!}(x-a)^2+\frac{f'''(a)}{3!}(x-a)^3+...\]

OpenStudy (amistre64):

so essentially, you want the first 5 derivatives of tan(x0

OpenStudy (experimentx):

if you wan't want to calculate that derivative ... there is a trick

OpenStudy (anonymous):

teach me that trick

OpenStudy (experimentx):

tan(x) = sin(x)/cos(x) let \[ \tan x = \sum_{k=0}^\infty a_kx^k \] then sin(x) = ??

OpenStudy (experimentx):

if you wan't want to calculate that *without* derivative ... there is a trick

OpenStudy (anonymous):

from the tanx = \[\sum_{k=o}^{\infty} akx ^{k}\] a is what

OpenStudy (experimentx):

\(a_k \) is the coefficient in the power series. we do not know it's value yet. do you know maclaurin series for sin(x) and cos(x)

OpenStudy (anonymous):

\[\sin x = \sum_{k=0}^{\infty} (-1)^k \frac{ x^{2k++!} }{ (2k+1)!} + 0 x^{2k+1} \] \[\cos x= \sum_{k=0}^{\infty} (-1)^K \frac{ x^2k }{ (2k)! } + 0 ( x^{2n} ) \]

OpenStudy (experimentx):

don't put that O notation if use use than infinity as your upper limit.

OpenStudy (experimentx):

\[ \sin x = \sum_{k=1}^\infty a_k x^k \times \cos x \\ \sum_{k=0}^\infty (-1)^k \frac{x^{2k+1}}{(2k+1)!} = \sum_{k=0}^\infty (-1)^k \frac{x^{2k}}{(2k)!} \times \sum_{k=0}^{\infty } a_k x^k \]

OpenStudy (amistre64):

just for the practice ... lets see if i can do this the "other" way ;) \[f:~tan\]\[f':~sec^2\]\[f'':~2~sec^3~sin\]\[f''':~6~sec^4~sin^2~+~2~sec^2\]\[f'''':~24~sec^5~sin^3~+~16~sec^3~sin\]\[f^{(5)}:~120~sec^6~sin^4~+~120~sec^4~sin^2~+~16~sec^2\]

OpenStudy (experimentx):

yeah you can do that too.

OpenStudy (experimentx):

Here's the trick \[ x-\frac{x^3}{6}+\frac{x^5}{120}-\frac{x^7}{5040}+O[x]^8 = \left( 1-\frac{x^2}{2}+\frac{x^4}{24}-\frac{x^6}{720}+O[x]^7 \right )\times \left( a_0 + a_1x + a_2 x^2 + a_3x^3 +...\right) \] up to five terms ... your job is to compare coefficients and determine their value. this will be yoru tan x

OpenStudy (amistre64):

the max series is at a=0 .... or simply evaluated at 0 f:0,2,4 all go to zero f:1,3,5 got to 1, 2, 16 \[P(x)=x+\frac{2}{3!}x^3+\frac{16}{5!}x^5+...\] i am definitely not smart enough to work thru experiments method ;)

OpenStudy (experimentx):

you are hard worker.

OpenStudy (experimentx):

\[ x-\frac{x^3}{6}+\frac{x^5}{120}-\frac{x^7}{5040}+O[x]^8 = \\ \left( 1-\frac{x^2}{2}+\frac{x^4}{24}-\frac{x^6}{720}+O[x]^7 \right )\times \left( a_0 + a_1x + a_2 x^2 + a_3x^3 +a_4x^4 +a_5x^5 + ...\right) \] there is no constant term on the left ... so \( a_0 *1 = 0 \) so \( a_0 = 0 \)

OpenStudy (experimentx):

on the left there is x ... from the right, you get \(a_1 x \) ... so a1 = 1

OpenStudy (experimentx):

keep doing it ... by watchig a2 and a4 = 0 and a3 = 1/3 a5 ... you have to determine by the combination of powers on the left. or just try hard way amistre mentioned.

OpenStudy (anonymous):

thank u very much... @experimentX and @amistre64 it very help me

OpenStudy (amistre64):

youre welcome :)